# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v2

@@ 2.2 Embodied Systems with Morphological Computation (System 0)
Biological systems are physical but also intelligent systems, i.e., their own bodies, outside of their cognitive systems. The system comprises pre-cognitive, embodied processes that occur at the super-fast timescale as our framework’s foundational layer of cognition. These processes emerge from direct physical interactions between an agent”s body and its environment without explicit computation, which is called morphological computation [104]. In other words, physical interactions generate the adaptive behavior of robots without computation using computers. Then, the system “computes” any tasks using the physical system. In addition, System 1 leverages System 0’s computing results for recognition, control, and learning.
@@
Morphological computation involves information processing by physical systems during the interaction between a robot’s body and the environment. The passive walker is a well-known example of morphological computation [88] (Figure 2a). This walker can walk down a slope without an actuator or computer. The physical interaction between the walker’s body, gravity, and slope generates the walking behavior. Another well-known example is a soft gripper [119]. This gripper can stably grasp objects of various shapes by passively changing its shape corresponding to the object and enclosing it with the same motor control signal.
@@
Morphological computation using physical interaction has been used for System 1’s control tasks of robots, as shown by the soft gripper. Iida et al. reported that a legged robot with elastic elements could stably maintain hopping by using its dynamics [63]. They showed that the elastic elements enable the hopping robot to maintain a stable attractor. Hamaya et al. proposed a method for learning peg insertion using a robotic arm with a soft wrist [48, 147] (Figure 2b). They divided peg insertion into multiple sub-tasks and moved the arm using simple commands in each sub-task. A robot using their method learns peg insertion efficiently owing to the simplicity of the commands in each sub-task and the passive deformation of the soft wrist corresponding to the contact state between the peg and hole. Preflex is a faster response of the human muscles than neural reflex [24]. Human runners adjust the stiffness of their ankles by co-activating antagonistic muscles to modulate the reaction force against the ground.
@@
Theories of physical reservoir computing (PRC) describe System 0’s computation using physical interaction [129]. PRC leverages the dynamical state changes of physical materials for computation, using sensor data as activation of nodes in RNNs. Nakajima et al. reported that the complex movement of an octopus-like soft arm can be used as a computational resource through PRC [96].
@@
Such computational results using physical interactions have been used for recognition. Homberg et al. developed a method to identify an object that is grasped by soft fingers [58]. The soft fingers change their shapes according to the shape of the grasped object. Therefore, the grasped object can be identified using the bending sensors on the fingers. Tanaka et al. reported that a flapping-wing aerial robot with soft wings and soft bending sensors on the wings could recognize the direction of the wind around the robot using PRC [130] (Figure 2c). In the flapping movements, the soft wings change their shape in a complex manner in response to the wind. The bending patterns of the wings reflect the direction of the wind. Thus, the direction can be classified using a simple linear classifier based on the time-series data of the bending sensors. Judd et al. proposed a method for estimating the location of an object in a water tank based on the movement of a soft arm in the tank using PRC [67]. The arm stirs the water, and the location of the object changes the water flow in the tank, which changes the movement of the arm. Therefore, the object’s location can be estimated from the time-series data of multiple marker positions on the arm.
@@ 2.3 Multi-agent Systems with Emergent Language
Humans have demonstrated remarkable capabilities as a species, largely attributed to the invention and use of language as a tool for collaboration. Language is believed to have co-evolved with the human brain [22], signifying a shift from individual to collective intelligence as the defining characteristic of our species. The idea of intelligence and life as collective phenomena has been a recurring topic of interest. The complexity and plasticity of language are fundamental to the formation of human societies and civilizations. Without these traits, the development of modern society and civilization would be difficult to imagine. In fields such as artificial intelligence, complex systems, and social decision-making, it is well established that collective intelligence often surpasses the capabilities of individual intelligence [127, 89]. In this study, we focused specifically on multi-agent systems (MASs) that utilize symbols and language.
@@
LLMs have demonstrated how statistically modeling language can form the basis of intelligence. However, although much of the research on LLMs emphasizes improving the performance of individual models, there is a growing trend of investigating LLMs as components of MASs. Recent studies have highlighted the potential of using multiple LLM-based agents in MASs to explore collective behaviors enabled by language. For example, Park et al., [100] presented a groundbreaking study titled "Generative Agents," where they used computational software agents based on LLMs to simulate the lives of 25 agents in a small virtual town. This study inspired subsequent research on utilizing LLM agents as models of human behavior, enabling rich multi-agent interactions that model human societies. Research in this area is rapidly expanding [41].
@@
Research has also demonstrated that leveraging multiple LLM agents can improve the overall system performance. For instance, Li et al., [82] showed that sampling and voting techniques, combined with increasing the number of agents, lead to enhanced LLM performance. Similarly, Wan et al., [150] proposed methods for knowledge fusion, where probabilistic distributions from multiple LLMs are aggregated to create a unified model. Mavromatis et al., [87] introduced a test-time fusion method, which uses perplexity-based approaches to integrate knowledge from multiple LLMs during generation.
@@
Another promising direction involves fostering collaboration through debates or discussions among agents to improve their reasoning capabilities. Chen et al., [19] proposed RECONCILE, a multi-model, multi-agent framework designed as a roundtable discussion among diverse LLM agents, significantly enhancing their reasoning capabilities. Motwani et al., [93] introduced Multi-Agent LLM Training (MALT), a training approach where multiple agents collaborate in a debate setting to solve reasoning tasks. Inspired by neural scaling laws, Qian et al., [108] explored how increasing the number of agents in multi-agent collaboration impacts performance. In addition, Subramaniam et al., [125] proposed the use of social interactions in multi-agent debates to fine-tune LLMs.
@@
While these approaches optimize the inference capabilities of MASs by leveraging pre-trained language models, another critical research avenue is to understand how such languages and symbols originally emerge. This area of study, known as emergent communication (EmCom), emergent language (EmLang), or symbol emergence, explores the dynamics of how languages evolve and adapt over time through decentralized human interactions [103, 80, 37, 124, 16, 136]. Language systems are inherently dynamic and change over time to describe a wide range of phenomena [22, 103, 121, 122, 153, 123, 144, 142]. Importantly, language is a type of symbol, and one of the most crucial characteristics of symbols is their arbitrariness. We argue that the plasticity of language and symbol systems is a key source of adaptability in humans.
@@
Throughout history, humans have adapted to their environments as collectives by transforming their symbolic systems, including their language. Building on discussions around EmCom, EmLang, and symbol emergence in robotics, Taniguchi, [136] introduced the collective predictive coding (CPC) hypothesis. This hypothesis reframes EmCom/EmLang from the perspective of generative cognitive systems. CPC posits that symbolic communication can be understood through generative modeling, both at the level of individual agents and across groups of agents. Based on the view of CPC, EmCom models have been developed as an extension of representation learning [47, 143, 36, 46, 64]. Beyond the emergence of external categorical signs and internal representations, Le Hoang et al., [81] explored the emergence of compositional word sequences within the framework of CPC, while You et al., [161] and Saito et al., [112] extended the framework to continuous signs, such as voice pitch and time-series data. For individual agents, the free energy principle (FEP) and active inference provide a generative framework for understanding cognitive processes. Extending this idea, CPC applies the principles of FEP and predictive coding (PC) at a societal scale [140], conceptualizing the emergence of symbols and language as a decentralized Bayesian inference process of shared external latent representations.
@@ 2.4 Bergson’s Philosophy: Multi-timescale interpretation
The temporal hierarchy related to life and mind is an important topic within philosophy as well. French philosopher Henri Bergson (1859–1941) is renowned for his distinctive theory of time. Recent studies, particularly from the Project Bergson in Japan (PBJ), have brought new insights into his work [54, 91, 134, 35, 160, 51, 52]. This new interdisciplinary approach is termed “Expanded Bergsonism” [90]. According to Hirai, [51], Bergson’s philosophy of time differs from traditional approaches by emphasizing multiple intrinsic timescales. This subsection outlines Bergson’s Multi-Timescale (MTS) interpretation in relation to System 0/1/2/3. The multi-timescale interpretation inspired by Bergson aligns naturally with contemporary cognitive architectures and robotics frameworks, particularly the hierarchical temporal dynamics observed in neuro-dynamical systems such as MTRNN and RNNPB, discussed in Section 2.1.2.
@@
Bergson’s concept of durée (duration), initially introduced to describe human conscious experience of time, was expanded in his second major book Matter and Memory [9] to encompass the temporality of matter and various life forms from the perspectives of emergence and evolution. This concept of “multiple rhythms of durées” can today be interpreted as a pluralism of timescales.
@@
Traditional philosophical treatments of time have generally viewed it either a single cosmic absolute time or as subjective psychological time, often placing them in dualistic opposition. The notion of multiple intrinsic timescales has largely been overlooked. In contrast, Bergson focused on the proper timescales inherent to each local interaction system. His renowned example of “sugar dissolving in water” illustrates that each event has its own duration, “which cannot be extended or shortened at will” ([10], p. 10). Different solutes or solvents require different times. Bergson’s philosophy resonates well with the concept of “characteristic time” in thermodynamics and condensed matter physics—concepts crucial to understanding system responses and relaxation processes.
@@
Importantly, cognitive activities themselves are natural phenomena that possesses intrinsic timescales. Just as heat diffusion or chemical reactions require a specific durations, organisms also need specific intervals to perceive situations and perform actions. Thus, sensorymotor processes can be viewed as “relaxation times” required to re-stabilize the system. Bergson defined the “thickness” of the present as follows: the real present “necessarily occupies a duration” and is “in its essence, sensorimotor” ([9], p. 153). While he did not specify precise values, contemporary scholars such as Wilhelm Wundt and recent studies suggest the “psychological present” spans approximately 3-5 seconds [102, 30, 154].
@@
Within this thick present, temporal consciousness (durée) resides. Changes that are too slow, such as celestial rotation or plant growth, are not directly experienced as “time perception.” Conversely, there is also a lower limit—the minimal temporal unit of perceptual elements. Bergson refers to “temporal resolution,” citing a value of around 2 millisecond ([9], p. 231, [29]). Changes occurring more rapidly, such as those in fluorescent lighting, similarly escape our perception of time, resulting in phenomena like the wagon-wheel effect.
@@
Thus, cognition operates within a specific temporal domain bounded by these upper and lower limits, termed here as the “window of the present.” Our consciousness perceives temporal changes exclusively within this domain. Bergson viewed this timescale-locality not as a limitation but as an enabling condition for consciousness itself. This particular “slowness,” relative to microscopic physical phenomena, is critical for the very emergence of consciousness [50, 51].
@@
Moreover, as Hirai ([53]) clarified, Bergson also described a dual-process model of cognition based on the flexible nature of the window of the present: “automatic recognition,” which remains confined to the thin present, and “attentive recognition,” which leverages the thickness of the present, roughly corresponding to modern Systems 1 and 2 (Bergson, 1896, chap. 2). This point will be revisited in Section 4.2.
@@
Bergson’s philosophy of time, particularly with its emphasis on multi intrinsic timescales, aligns with certain aspects of dual-process theories. Furthermore, it offers potential for broader discussions concerning the philosophy of consciousness and more expansive conceptions of time. Extending dual-process theory and elucidating its connections to Bergson’s philosophy of time could prove crucial for advancing philosophical discourse in these areas.
@@ 3.1 System 0/1/2/3
For Systems 1 and 2, we follow the conventional definitions [69]. In connection with Bergson’s philosophy, note that the dual temporal structure of cognition described in Section 2.1 aligns closely with Bergson’s dual-memory model—automatic recognition utilizing motor memory and attentive recognition leveraging pure memory (see Section 4.2).
@@
System 3 represents the highest level of (super-)cognitive processing in our proposed framework, encompassing collective intelligence and symbol emergence. If the CPC hypothesis is correct, we as language creators engage in predictive coding or free energy minimization as an entire system [136].

### v3

@@ 2.2 Embodied Systems with Morphological Computation (System 0)
Biological systems are physical but also intelligent systems, i.e., their own bodies, outside of their cognitive systems. The system comprises pre-cognitive, embodied processes that occur at the super-fast timescale as our framework’s foundational layer of cognition. These processes emerge from direct physical interactions between an agent”s body and its environment without explicit computation, which is called morphological computation (pfeifer2001understanding). In other words, physical interactions generate the adaptive behavior of robots without computation using computers. Then, the system “computes” any tasks using the physical system. In addition, System 1 leverages System 0’s computing results for recognition, control, and learning.
@@
Morphological computation involves information processing by physical systems during the interaction between a robot’s body and the environment. The passive walker is a well-known example of morphological computation (mcgeer1990passive) (Figure 2a). This walker can walk down a slope without an actuator or computer. The physical interaction between the walker’s body, gravity, and slope generates the walking behavior. Another well-known example is a soft gripper (shintake2018soft). This gripper can stably grasp objects of various shapes by passively changing its shape corresponding to the object and enclosing it with the same motor control signal.
@@
Morphological computation using physical interaction has been used for System 1’s control tasks of robots, as shown by the soft gripper. Iida et al. reported that a legged robot with elastic elements could stably maintain hopping by using its dynamics (iida2004cheap). They showed that the elastic elements enable the hopping robot to maintain a stable attractor. Hamaya et al. proposed a method for learning peg insertion using a robotic arm with a soft wrist (hamaya2020learning; von2020compact) (Figure 2b). They divided peg insertion into multiple sub-tasks and moved the arm using simple commands in each sub-task. A robot using their method learns peg insertion efficiently owing to the simplicity of the commands in each sub-task and the passive deformation of the soft wrist corresponding to the contact state between the peg and hole. Preflex is a faster response of the human muscles than neural reflex (dickinson2000animals). Human runners adjust the stiffness of their ankles by co-activating antagonistic muscles to modulate the reaction force against the ground.
@@
Theories of physical reservoir computing (PRC) describe System 0’s computation using physical interaction (tanaka2019recent). PRC leverages the dynamical state changes of physical materials for computation, using sensor data as activation of nodes in RNNs. Nakajima et al. reported that the complex movement of an octopus-like soft arm can be used as a computational resource through PRC (nakajima2014exploiting).
@@
Such computational results using physical interactions have been used for recognition. Homberg et al. developed a method to identify an object that is grasped by soft fingers (homberg2015haptic). The soft fingers change their shapes according to the shape of the grasped object. Therefore, the grasped object can be identified using the bending sensors on the fingers. Tanaka et al. reported that a flapping-wing aerial robot with soft wings and soft bending sensors on the wings could recognize the direction of the wind around the robot using PRC (tanaka2021flapping) (Figure 2c). In the flapping movements, the soft wings change their shape in a complex manner in response to the wind. The bending patterns of the wings reflect the direction of the wind. Thus, the direction can be classified using a simple linear classifier based on the time-series data of the bending sensors. Judd et al. proposed a method for estimating the location of an object in a water tank based on the movement of a soft arm in the tank using PRC (judd2019sensing). The arm stirs the water, and the location of the object changes the water flow in the tank, which changes the movement of the arm. Therefore, the object’s location can be estimated from the time-series data of multiple marker positions on the arm.
@@ 2.3 Multi-agent Systems with Emergent Language
Humans have demonstrated remarkable capabilities as a species, largely attributed to the invention and use of language as a tool for collaboration. Language is believed to have co-evolved with the human brain (deacon1998symbolic), signifying a shift from individual to collective intelligence as the defining characteristic of our species. The idea of intelligence and life as collective phenomena has been a recurring topic of interest. The complexity and plasticity of language are fundamental to the formation of human societies and civilizations. Without these traits, the development of modern society and civilization would be difficult to imagine. In fields such as artificial intelligence, complex systems, and social decision-making, it is well established that collective intelligence often surpasses the capabilities of individual intelligence (surowiecki2004wisdom; miller2010smart). In this study, we focused specifically on multi-agent systems (MASs) that utilize symbols and language.
@@
LLMs have demonstrated how statistically modeling language can form the basis of intelligence. However, although much of the research on LLMs emphasizes improving the performance of individual models, there is a growing trend of investigating LLMs as components of MASs. Recent studies have highlighted the potential of using multiple LLM-based agents in MASs to explore collective behaviors enabled by language. For example, park2023generative presented a groundbreaking study titled "Generative Agents," where they used computational software agents based on LLMs to simulate the lives of 25 agents in a small virtual town. This study inspired subsequent research on utilizing LLM agents as models of human behavior, enabling rich multi-agent interactions that model human societies. Research in this area is rapidly expanding (guo2024large).
@@
Research has also demonstrated that leveraging multiple LLM agents can improve the overall system performance. For instance, li2024more showed that sampling and voting techniques, combined with increasing the number of agents, lead to enhanced LLM performance. Similarly, wan2024knowledge proposed methods for knowledge fusion, where probabilistic distributions from multiple LLMs are aggregated to create a unified model. mavromatis2024pack introduced a test-time fusion method, which uses perplexity-based approaches to integrate knowledge from multiple LLMs during generation.
@@
Another promising direction involves fostering collaboration through debates or discussions among agents to improve their reasoning capabilities. chen2023reconcile proposed RECONCILE, a multi-model, multi-agent framework designed as a roundtable discussion among diverse LLM agents, significantly enhancing their reasoning capabilities. motwani2024malt introduced Multi-Agent LLM Training (MALT), a training approach where multiple agents collaborate in a debate setting to solve reasoning tasks. Inspired by neural scaling laws, qian2024scaling explored how increasing the number of agents in multi-agent collaboration impacts performance. In addition, subramaniam2025multiagent proposed the use of social interactions in multi-agent debates to fine-tune LLMs.
@@
While these approaches optimize the inference capabilities of MASs by leveraging pre-trained language models, another critical research avenue is to understand how such languages and symbols originally emerge. This area of study, known as emergent communication (EmCom), emergent language (EmLang), or symbol emergence, explores the dynamics of how languages evolve and adapt over time through decentralized human interactions (peters2024survey; LazaridouB-2020-emergent; galke2025learning; steels2015talking; cangelosi1998emergence; taniguchi2024collective). Language systems are inherently dynamic and change over time to describe a wide range of phenomena (deacon1998symbolic; peters2024survey; Steels97; Steels2007; wittgenstein2009philosophical; STEELS2011339; tomasello2005constructing; taniguchi2018symbol). Importantly, language is a type of symbol, and one of the most crucial characteristics of symbols is their arbitrariness. We argue that the plasticity of language and symbol systems is a key source of adaptability in humans.
@@
Throughout history, humans have adapted to their environments as collectives by transforming their symbolic systems, including their language. Building on discussions around EmCom, EmLang, and symbol emergence in robotics, taniguchi2024collective introduced the collective predictive coding (CPC) hypothesis. This hypothesis reframes EmCom/EmLang from the perspective of generative cognitive systems. CPC posits that symbolic communication can be understood through generative modeling, both at the level of individual agents and across groups of agents. Based on the view of CPC, EmCom models have been developed as an extension of representation learning (hagiwara2019symbol; taniguchi2022emergent; furukawa2022symbol; hagiwara2022multiagent; inukai2023recursive). Beyond the emergence of external categorical signs and internal representations, hoang2024compositionality explored the emergence of compositional word sequences within the framework of CPC, while you2024multimodal and saito2024emergence extended the framework to continuous signs, such as voice pitch and time-series data. For individual agents, the free energy principle (FEP) and active inference provide a generative framework for understanding cognitive processes. Extending this idea, CPC applies the principles of FEP and predictive coding (PC) at a societal scale (taniguchi2024cpcms), conceptualizing the emergence of symbols and language as a decentralized Bayesian inference process of shared external latent representations.
@@ 2.4 Bergson’s Philosophy: Multi-timescale interpretation
The temporal hierarchy related to life and mind is an important topic within philosophy as well. French philosopher Henri Bergson (1859–1941) is renowned for his distinctive theory of time. Recent studies, particularly from the Project Bergson in Japan (PBJ), have brought new insights into his work (hirai2016; miyake_hirai_2018; tani2018; fujita2022; yoneda2022; hirai2022; hirai2023). This new interdisciplinary approach is termed “Expanded Bergsonism” (miquel2023). According to hirai2022, Bergson’s philosophy of time differs from traditional approaches by emphasizing multiple intrinsic timescales. This subsection outlines Bergson’s Multi-Timescale (MTS) interpretation in relation to System 0/1/2/3. The multi-timescale interpretation inspired by Bergson aligns naturally with contemporary cognitive architectures and robotics frameworks, particularly the hierarchical temporal dynamics observed in neuro-dynamical systems such as MTRNN and RNNPB, discussed in Section 2.1.2.
@@
Bergson’s concept of durée (duration), initially introduced to describe human conscious experience of time, was expanded in his second major book Matter and Memory (bergson1896) to encompass the temporality of matter and various life forms from the perspectives of emergence and evolution. This concept of “multiple rhythms of durées” can today be interpreted as a pluralism of timescales.
@@
Traditional philosophical treatments of time have generally viewed it either a single cosmic absolute time or as subjective psychological time, often placing them in dualistic opposition. The notion of multiple intrinsic timescales has largely been overlooked. In contrast, Bergson focused on the proper timescales inherent to each local interaction system. His renowned example of “sugar dissolving in water” illustrates that each event has its own duration, “which cannot be extended or shortened at will” (bergson1907, p. 10). Different solutes or solvents require different times. Bergson’s philosophy resonates well with the concept of “characteristic time” in thermodynamics and condensed matter physics—concepts crucial to understanding system responses and relaxation processes.
@@
Importantly, cognitive activities themselves are natural phenomena that possesses intrinsic timescales. Just as heat diffusion or chemical reactions require a specific durations, organisms also need specific intervals to perceive situations and perform actions. Thus, sensorymotor processes can be viewed as “relaxation times” required to re-stabilize the system. Bergson defined the “thickness” of the present as follows: the real present “necessarily occupies a duration” and is “in its essence, sensorimotor” (bergson1896, p. 153). While he did not specify precise values, contemporary scholars such as Wilhelm Wundt and recent studies suggest the “psychological present” spans approximately 3-5 seconds (poppel1985; fraisse1957; wittmann2016felt).
@@
Within this thick present, temporal consciousness (durée) resides. Changes that are too slow, such as celestial rotation or plant growth, are not directly experienced as “time perception.” Conversely, there is also a lower limit—the minimal temporal unit of perceptual elements. Bergson refers to “temporal resolution,” citing a value of around 2 millisecond (bergson1896, p. 231, exner1894). Changes occurring more rapidly, such as those in fluorescent lighting, similarly escape our perception of time, resulting in phenomena like the wagon-wheel effect.
@@
Thus, cognition operates within a specific temporal domain bounded by these upper and lower limits, termed here as the “window of the present.” Our consciousness perceives temporal changes exclusively within this domain. Bergson viewed this timescale-locality not as a limitation but as an enabling condition for consciousness itself. This particular “slowness,” relative to microscopic physical phenomena, is critical for the very emergence of consciousness (hirai2019; hirai2022).
@@
Moreover, as hirai2023b clarified, Bergson also described a dual-process model of cognition based on the flexible nature of the window of the present: “automatic recognition,” which remains confined to the thin present, and “attentive recognition,” which leverages the thickness of the present, roughly corresponding to modern Systems 1 and 2 (Bergson, 1896, chap. 2). This point will be revisited in Section 4.2.
@@
Bergson’s philosophy of time, particularly with its emphasis on multiple intrinsic timescales, aligns with certain aspects of dual-process theories. Furthermore, it offers potential for broader discussions concerning the philosophy of consciousness and more expansive conceptions of time. Extending dual-process theory and elucidating its connections to Bergson’s philosophy of time could prove crucial for advancing philosophical discourse in these areas.
@@ 3.1 System 0/1/2/3
For Systems 1 and 2, we follow the conventional definitions (kahneman2011thinking). In connection with Bergson’s philosophy, note that the dual temporal structure of cognition described in Section 2.1 aligns closely with Bergson’s dual-memory model—automatic recognition utilizing motor memory and attentive recognition leveraging pure memory (see Section 4.2).
@@
System 3 represents the highest level of (super-)cognitive processing in our proposed framework, encompassing collective intelligence and symbol emergence. If the CPC hypothesis is correct, we as language creators engage in predictive coding or free energy minimization as an entire system (taniguchi2024collective).